# Understanding Blockchain Governance: Analyzing Decentralized Voting to Amend DeFi Smart Contracts

### v2

@@ 4.2. Voting on Governance Proposals
To propose changes to the Compound protocol, an address must have at least $25000$ COMP tokens delegated to it to create a proposal. [^4] Prior proposal $89$, an address should have at least $65000$ delegated tokens to create proposals. However, as of Sept. 18, 2021, proposal #60 introduced an exception to this rule, allowing also whitelisted-addresses to create proposals even if they do not have $25000$ delegated tokens.
@@
In total, $3335$ Compound voters cast their votes through $9500$ transactions with $8769$ ($92.31$%) for in-favor votes, $644$ ($6.78$%) for against votes, and finally $87$ ($0.91$%) for abstained votes. Most of the voters ($51.36$%) only voted for $1$ proposal. $1$% of participants voted for at least $26.66$ proposals. On average, participants voted on $2.85$ proposals with a standard deviation (std.) of $5.23$. The address 0x84e3normal-$\cdots$5a95 voted on the maximum number of proposals ($100$), followed by MonetSupply and blck who voted on $96$ and $88$ proposals, respectively.
Figure 4. The lifecycle of a Compound proposal lasts 7 days. After a proposal is created, it waits for 2 days before the 3-day voting period begins. Once the outcome of the election is decided, it takes 2 more days for the proposal to be executed and become part of the Compound Governance protocol. Proposals can also be cancelled at any time before they are executed.
Creation of proposals  In total, $33$ Compound proposers created the 133 proposals. Of these proposers, $16$ ($48.48$%) created one proposal, while $10$% of them created at least $8$ proposals. The average number of proposals created per proposer is $4.03$ proposals, with a std. of $5.27$ and a median of $2$. The highest number of proposals was created by Gauntlet, who created $24$ proposals, followed by blck, who created $20$ proposals.
The maximum number of proposals were created in Mar. 2022, $11$ proposals created (from #86 to #96). However, of those, only $5$ were executed, as $1$ was defeated and $5$ were cancelled (see Fig. 5). Proposals were submitted, on average, every $6.95$ days (std. of $6.41$), with a median of $5.08$ days. The shortest and longest interval between proposals was 00 and $31.14$ days, respectively. Proposals typically take $1.64$ days (std. of $0.72$) to reach the quorum (Fig. 19 in §H).
@@
Figure 5. Monthly number of Compound proposals created and their outcome (executed, defeated, and cancelled).
Participation in voting  Next, we computed the voting participation per proposal (see Fig. 6). This metric is calculated by dividing the number of votes (or delegated tokens) cast on a proposal by the total number of delegated tokens eligible to vote on that proposal at the start of the voting period. This is a crucial measurement as it shows the proportion of all delegated tokens that are used in the governance election process by the voters on proposals. Also, protocols with low voter turnout are more susceptible to vote-buying, as non-voting users may sell their voting rights to others (Daian et al., 2018). Our results show that the average voter turnout is $33.25$% (std. of $17.61$%), the median is $32.10$%, and the maximum $80.80$%. Based on Fig. 6, we observe higher voting participation for early proposals compared to recent ones, likely due to the limited availability of tokens to a select few in the beginning.
@@
Figure 6. Compound’s voting participation per proposal in terms of delegated tokens used from all delegated tokens available. Proposals are indicated either as executed (in green), defeated (in red), or cancelled (in blue).
@@
Furthermore, we analyzed the number of voters needed for proposals to reach $50$% of the total votes cast. Out of $133$ proposals, we excluded $7$ proposals that were cancelled before the voting period, leaving us with $126$ ($94.74$%) proposals for analysis. On average, those proposals required $2.84$ voters with a std. of $0.97$ and a median of $3$ voters. The minimum and maximum number of voters were $1$ and $5$, respectively. This again suggests that the token distribution is concentrated among few voters who hold a high voting power. Fig. 20 in §G shows the cumulative voting power for the top 10 most powerful voters for each of these $126$ proposals in our data set.
@@
Margin of victory/defeat  During the analyzed period, $133$ proposals have been created. Of these, $17$ were cancelled and $15$ defeated, leaving $101$ ($75.94$%) executed proposals. Fig. 10 shows the percentage of in-favor, against, and abstain votes for Compound each proposal. The majority of the proposals received significant support from the voters. On average, proposals received $89.39$% of the votes in favor, with a std. of $23.98$% votes and a median of $99.99$%. We highlight the proposals’ outcome at each stage of their lifecycle in Fig. 11. Our analyses show that $7$ ($5.26$%) out of $133$ proposals were cancelled right after they were created and, therefore, they had not reached the Voting Period meaning they were not available for voting. Next, 4 proposals were cancelled before the Voting Ends stage, meaning they were pulled out before the election finished. $2$ were also cancelled after they succeeded in the election (after the Voting Ends stage) but before they were queued in the Timelock. Further, $4$ proposals were cancelled when in the Timelock. These proposals account for $6$ cancelled proposals after they successfully passed, which could indicate a lack of community consensus (Sharma et al., 2023). Finally, $101$ ($75.94$)proposals were successfully executed. We report the results for Uniswap in Fig. 16 in  §A. Further, we gathered data from Messari (Messari, 2023) to categorize the executed Compound proposals and report their importance level in  §B. In §C we also discuss the temporal dynamics of voting.
@@ 4.3. Voting patterns of delegates
In this section, we analyze the formation of coalitions among voters, where they cast their votes as a group. This analysis is crucial because such behavior may compromise the security of the governance protocol. Specifically, instead of expressing their individual opinions on a proposal, voters may choose to mimic the votes of their peers. The transparency of the Ethereum blockchain used for voting in Compound allows anyone to view the addresses of voters and their corresponding votes (e.g., their voting power and voting preference) during the election process, potentially facilitating this behavior. As a result, exploring the possibility of coalition formation could provide valuable insights into the decision-making patterns of voters. Fig. 12 shows a heatmap of how each of the top $15$ voters cast their votes across all $133$ proposals in our data set, where in-favor votes are in green, against in red, and abstain in blue.
@@
5. Related work
Decentralized Autonomous Organizations (DAOs)  Several prior studies analyzed the governance structures of DAOs (Beck et al., 2018; Rikken et al., 2019; Hassan and De Filippi, 2021). Hassan and De Filippi analyzed what DAOs constitute and discuss their key traits (Hassan and De Filippi, 2021). Rikken et al. identified various political challenges in governance of blockchains (Rikken et al., 2019). Beck et al. (Beck et al., 2018) presented a case study of a DAO in Swarm City (City, 2023), a decentralized commerce platform. A recent work categorized the governance of several blockchains such as Bitcoin, Ethereum, Tezos, Polkadot, and some governance protocols like Uniswap (Adams et al., 2021), MakerDAO (MakerDAO, 2023) and Compound (Leshner and Hayes, 2019) into different types (Kiayias and Lazos, 2023). These invaluable prior work do not, nevertheless, empirically examine the data on existing DAOs to characterize how users interact with on-chain governance smart contracts.
6. Concluding discussion
We conduct in-dept analyses of Compound and Uniswap, two widely used smart contracts. Our analysis centers on their decentralized governance protocols, with a particular focus on amendments to their contract. We found that the Compound contract is being actively amended—token holders continuously propose amendments that are then voted on by other token holders. We observed a striking concentration of tokens in the hands of a few participants for both Compound and Uniswap, which raises serious concerns about the extent to which governance is decentralized in practice. For instance, our analysis showed that, on average, only $3.25$ voters were needed for the proposals to reach quorum and pass in Compound, and only $2.84$ voters were needed to reach $50$% of the total votes. Our analysis also highlights issues with the Compound use of on-chain voting – in particular, the transaction fees voters must pay to cast an on-chain vote can make it prohibitively expensive for voters with fewer tokens. These costs have implications for voting participation and can affect how voters, proposers, and other stakeholders interact with these protocols.
Our findings motivate new research directions; for example, we find empirical evidence in support of recent proposals to redefine voting power based on social rewards, such as a voter’s reputation or a voter’s contributions to the protocol (Sharma et al., 2023; Liu et al., 2022; Guidi et al., 2021) or the use of a quadratic voting scheme, where voting power is calculated as the square root of the number of tokens held by the voters (Buterin et al., 2019; Lalley and Weyl, 2018).
Acknowledgments
Appendix A Analysis of Uniswap’s Governance
@@
We gathered transaction and events data from our Ethereum archive node regarding Uniswap since its conception. Tab. 1 and Tab. 2 show the data set description we gathered from our node that was used in the analysis for the Uniswap (UNI) token and the Uniswap Governance protocol, respectively.
@@
(a) Tokens in circulation
(b) Tokens delegated
Figure 14. Distribution of the top 15 COMP tokens per accounts on Nov. 7, 2022: (a) These accounts hold 33.27% (187.61 million) out of 563.92 million UNI tokens in circulation; (b) These addresses have 56.37% (117.27 million) of all delegated tokens.
Proposal Creation and voting cast  Uniswap proposals were submitted, on average, every $16.81$ days (std. of $22.62$), with a median of $5.52$ days, as seen in Fig. 14(a). The shortest and longest interval between proposals was 00 and $80.98$ days, respectively. During August and September 2021, the largest number of proposals, totaling 10, were generated. It is noteworthy that a significant proportion of these proposals were initiated by the Uniswap Governor as initial proposals, with support levels ranging from 2.1 to 2.8. These proposals did not receive any votes and, are also not documented on the Uniswap webpage (Labs, 2023c). Unfortunately, we were unable to verify the reasons that could explain this.
@@
Appendix B Proposals Categorization
We gathered data from Messari (Messari, 2023) to determine the categories, subcategories, and the level of importance associated with each Compound proposal. Fig. 17 shows the distribution of $101$ executed Compound proposals across different categories and subcategories. We show the degree of importance for each proposal according to Messari divided into “low”, “medium”, “high”, and “very high”. As a result, a few proposals categorized as “Parameter Change” and “Security” demonstrate a high level of importance. Furthermore, proposals with the highest level of importance are found within the “Security” category, specifically within the “Mining and Validation” subcategory. This refers to the proposal 64 that was created to fix a bug introduced by proposal #62.
Appendix C Temporal dynamics of voting
@@
According to our analysis, voters take an average of $1.4$ days (with a std. of $0.95$ and a median of $1.34$ days) to cast their votes after the voting period began. The shortest and longest recorded delays in our data set are 00 and $3.39$ days, respectively. Fig. 18 shows the distribution of the time it takes voters to cast their votes for each proposal. We also highlight the voting delays for all votes cast per proposal in Fig. 21 in §G.
@@
Appendix D Filtering events to construct our Compound data set
@@
Protocol Type Voting Who can vote? Delegation Voting Aggregation How proposals are implemented
AAVE (AAVE, 2023) Lending on-chain
addresses with delegated tokens
yes on-chain
on-chain via an SC call.
Balancer (Balancer.fi, 2023) DEX off-chain
stakers with locked tokens
yes (off-chain) off-chain
via 6-of-11 multisig.
Compound (Leshner and Hayes, 2019) Lending on-chain
addresses with delegated tokens
yes on-chain
on-chain via an SC call.
Convex Finance (Convex, 2023a) Yield Farming off-chain
stakers with locked tokens
yes (off-chain) off-chain
via 3-of-5 multisig.
Curve (Curve, 2023) DEX on-chain
holders
yes on-chain
on-chain through an SC call.
Maker Executive (MakerDAO, 2023) Stablecoin on-chain
holders
no on-chain
New Governance Contract requires more MKR staked than previous.
Maker Polling (MakerDAO, 2023) Stablecoin on-chain
addresses with delegated tokens
yes off-chain
Engineers at Maker create the governance contract based on the voting outcome.
Uniswap (Adams et al., 2021) DEX on-chain
addresses with delegated tokens
yes on-chain
on-chain via an SC call.
@@
Appendix E Inferring Wallet Addresses Ownership
@@
Furthermore, to extend the available identified addresses in our analysis, we obtained the addresses of $2783$ identified users from the Sybil-List (Sybil, 2023b), a project maintained by Uniswap that uses cryptographic proofs to verify wallet addresses ownership. By combining the identified addresses from both sources, we were able to obtain the ownership of $3191$ inferred public wallet addresses to use in our analysis. We were able to infer $114$ ($3.41$%) out of the $3341$ unique addresses in our data set. Considering the top 10 most powerful voters for each proposal (refer to Fig. 20 in §G), we were able to infer $67$ ($50.37$%) of the $133$ unique addresses. Overall, our methodology allowed us to partially overcome the anonymity of public wallet addresses on the Ethereum blockchain and shed light on the ownership of these addresses in our data set. Finally, as an entity can control more than one address, we grouped the addresses we identified belonging to the same entity together to conduct our analysis.
Appendix F Types of Existing Governance Protocols
There are various smart contract applications that utilize decentralized governance protocols for decision-making, including those for lending, decentralized exchanges (DEXes), and stablecoins, among others. An example of such protocols can be found on the Ethereum blockchain, where a number of these applications are available. We have selected some of the most protocols that use decentralized governance for decision-making. Tab. 3 presents 8 protocols, including Maker Executive and Maker Pooling, which are part of the MakerDAO (MakerDAO, 2023) stablecoin protocol responsible for the DAI token. These protocols use decentralized governance mechanisms, and we characterize them based on whether their votes are cast on- or off-chain, the delegation methods they use, how they aggregate the votes, and how the proposal outcome take effect.
Appendix G How voters cast their votes
This section examines how each of the top-10 voters of Compound and Uniswap cast their votes. Some proposals may not have received any votes if they were cancelled before the voting period began. See §10 for details. Fig. 20 shows how each of the top-10 voters cast their votes in each of the $126$ ($94.74$%) out of $133$ Compound proposals.

### v3

@@ 4.2. Voting on Governance Proposals
To propose changes to the Compound protocol, an address must have at least $25000$ COMP tokens delegated to it to create a proposal. [^4] Prior proposal $89$, an address should have at least $65000$ delegated tokens to create proposals (at Berkeley, 2021). However, as of Sept. 18, 2021, proposal #60 introduced an exception to this rule, allowing also whitelisted-addresses to create proposals even if they do not have $25000$ delegated tokens (Capital, 2021).
@@
In total, $3335$ Compound voters cast their votes through $9500$ transactions with $8769$ ($92.31$%) for in-favor votes, $644$ ($6.78$%) for against votes, and finally $87$ ($0.91$%) for abstained votes. Most of the voters ($51.36$%) only voted for $1$ proposal. $1$% of participants voted for at least $26.66$ proposals. On average, participants voted on $2.85$ proposals with a standard deviation (std.) of $5.23$. The address 0x84e3$\cdots$5a95 voted on the maximum number of proposals ($100$), followed by MonetSupply and blck who voted on $96$ and $88$ proposals, respectively.
Figure 4. The lifecycle of a Compound proposal until it gets executed lasts 7 days.
4.2.1. Creation of proposals
In total, $33$ Compound proposers created the 133 proposals. Of these proposers, $16$ ($48.48$%) created one proposal, while $10$% of them created at least $8$ proposals. The average number of proposals created per proposer is $4.03$ proposals, with a std. of $5.27$ and a median of $2$. The highest number of proposals was created by Gauntlet, who created $24$ proposals, followed by blck, who created $20$ proposals.
The maximum number of proposals were created in Mar. 2022, $11$ proposals created (from #86 to #96). However, of those, only $5$ were executed, as $1$ was defeated and $5$ were cancelled (see Fig. 5). Proposals were submitted, on average, every $6.95$ days (std. of $6.41$), with a median of $5.08$ days. The shortest and longest interval between proposals was 00 and $31.14$ days, respectively. Proposals typically take $1.64$ days (std. of $0.72$) to reach the quorum (Fig. 19 in §E).
@@ 4.2.1. Creation of proposals
Figure 5. Monthly number of Compound proposals created and their final outcome.
4.2.2. Participation in voting
Next, we computed the voting participation per proposal (see Fig. 6). This metric is calculated by dividing the number of votes (or delegated tokens) cast on a proposal by the total number of delegated tokens eligible to vote on that proposal at the start of the voting period. This is a crucial measurement as it shows the proportion of all delegated tokens that are used in the governance election process by the voters on proposals. Also, protocols with low voter turnout are more susceptible to vote-buying, as non-voting users may sell their voting rights to others (Daian et al., 2018). Our results show that the average voter turnout is $33.25$% (std. of $17.61$%), the median is $32.10$%, and the maximum $80.80$%. Based on Fig. 6, we observe higher voting participation for early proposals compared to recent ones, likely due to the limited availability of tokens to a select few in the beginning.
@@ 4.2.2. Participation in voting
Figure 6. Compound’s voting participation per proposal in terms of delegated tokens used from all delegated tokens available.
@@
Furthermore, we analyzed the number of voters needed for proposals to reach $50$% of the total votes cast. Out of $133$ proposals, we excluded $7$ proposals that were cancelled before the voting period, leaving us with $126$ ($94.74$%) proposals for analysis. On average, those proposals required $2.84$ voters with a std. of $0.97$ and a median of $3$ voters. The minimum and maximum number of voters were $1$ and $5$, respectively. This again suggests that the token distribution is concentrated among few voters who hold a high voting power. Fig. 20 in §D shows the cumulative voting power for the top 10 most powerful voters for each of these $126$ proposals in our data set.
@@
4.2.3. Margin of victory/defeat
During the analyzed period, $133$ proposals have been created. Of these, $17$ were cancelled and $15$ defeated, leaving $101$ ($75.94$%) executed proposals. Fig. 10 shows the percentage of in-favor, against, and abstain votes for Compound each proposal. The majority of the proposals received significant support from the voters. On average, proposals received $89.39$% of the votes in favor, with a std. of $23.98$% votes and a median of $99.99$%. We highlight the proposals’ outcome at each stage of their lifecycle in Fig. 11. Our analyses show that $7$ ($5.26$%) out of $133$ proposals were cancelled right after they were created and, therefore, they had not reached the Voting Period meaning they were not available for voting. Next, 4 proposals were cancelled before the Voting Ends stage, meaning they were pulled out before the election finished. $2$ were also cancelled after they succeeded in the election (after the Voting Ends stage) but before they were queued in the Timelock. Further, $4$ proposals were cancelled when in the Timelock. These proposals account for $6$ cancelled proposals after they successfully passed, which could indicate a lack of community consensus (Sharma et al., 2023). Finally, $101$ ($75.94$)proposals were successfully executed. We report the results for Uniswap in Fig. 17 in  §5. Further, we gathered data from Messari (Messari, 2023) to categorize the executed Compound proposals and report their importance level in  §4.3. In §A we also discuss the temporal dynamics of voting.
@@ 4.2.3. Margin of victory/defeat
4.3. Proposals Categorization
In this section, we explore the importance of each type of proposal that have been implemented, examining their impact on the Compound protocol. Specifically, we explore whether these proposals mainly focus on adjusting parameters or addressing security concerns.
To this end, we gathered data from Messari (Messari, 2023) to determine the categories, subcategories, and the level of importance associated with each Compound proposal. Fig. 12 shows the distribution of $101$ executed Compound proposals across different categories and subcategories. We show the degree of importance for each proposal according to Messari divided into “low”, “medium”, “high”, and “very high”. As a result, a few proposals categorized as “Parameter Change” and “Security” demonstrate a high level of importance. Furthermore, proposals with the highest level of importance are found within the “Security” category, specifically within the “Mining and Validation” subcategory. This refers to the proposal 64 that was created to fix a bug introduced by proposal #62.
@@ 4.4. Voting patterns of delegates
In this section, we analyze the formation of coalitions among voters, where they cast their votes as a group. This analysis is crucial because such behavior may compromise the security of the governance protocol. Specifically, instead of expressing their individual opinions on a proposal, voters may choose to mimic the votes of their peers. The transparency of the Ethereum blockchain used for voting in Compound allows anyone to view the addresses of voters and their corresponding votes (e.g., their voting power and voting preference) during the election process, potentially facilitating this behavior. As a result, exploring the possibility of coalition formation could provide valuable insights into the decision-making patterns of voters. Fig. 13 shows a heatmap of how each of the top $15$ voters cast their votes across all $133$ proposals in our data set.
@@
5. An Analysis of Uniswap’s Governance
In this section, we provide more details into the Uniswap and its governance protocol. Uniswap (Adams et al., 2021) is a decentralized exchange (DEX) platform that enables users to trade one token for another. In addition, users can add liquidity to Uniswap and earn interest based on the amount they provide.
The Uniswap token (UNI) is an ERC-20 token that allows its holders to transfer UNI between addresses. Similarly to Compound, with Uniswap, holders can also participate in the Uniswap Governance Protocol by proposing changes and voting on proposals. Uniswap Governance is derived from Compound Governance Alpha and Bravo (refer to §4), and therefore shares similar features. For example, token holders must delegate their tokens to themselves or other addresses in order to vote. Each delegated token counts as one vote. Similar to Compound, participants can increase their voting power by holding multiple tokens.
@@ 5. An Analysis of Uniswap’s Governance
We gathered transaction and events data from our Ethereum archive node regarding Uniswap since its conception. Table 2 and Table 3 show the data set description we gathered from our node that was used in the analysis for the Uniswap (UNI) token and the Uniswap Governance protocol, respectively.
@@
(a) UNI tokens in circulation
(b) UNI tokens delegated
Figure 15. Distribution of the top 15 UNI tokens per accounts on Nov. 7, 2022: (a) These accounts hold 33.27% (187.61 million) out of 563.92 million UNI tokens in circulation; (b) These addresses have 56.37% (117.27 million) of all delegated tokens.
Proposal Creation and voting cast  Uniswap proposals were submitted, on average, every $16.81$ days (std. of $22.62$), with a median of $5.52$ days, as seen in Fig. 16(a). The shortest and longest interval between proposals was 00 and $80.98$ days, respectively. During August and September 2021, the largest number of proposals, totaling 10, were generated. It is noteworthy that a significant proportion of these proposals were initiated by the Uniswap Governor as initial proposals, with support levels ranging from 2.1 to 2.8. These proposals did not receive any votes and, are also not documented on the Uniswap webpage (Labs, 2023c). Upon conducting a further analysis, it was observed that the “target” field within their proposals has an Ethereum null address (i.e., 0x0). This means that, even in the event of these proposals passing, no modifications would be introduced to any contracts. The inclusion of these proposals seems to be a preemptive measure, possibly aimed at preventing conflicts with proposal numbers in the Governor Alpha contract. However, we could not verify that.
@@
6. Related work
Decentralized Autonomous Organizations (DAOs)  Several prior studies analyzed the governance structures of DAOs (Beck et al., 2018; Rikken et al., 2019; Hassan and De Filippi, 2021; Dotan et al., 2023; Tan et al., 2023; Austgen et al., 2023). Hassan and De Filippi analyzed what DAOs constitute and discuss their key traits (Hassan and De Filippi, 2021). Rikken et al. identified various political challenges in governance of blockchains (Rikken et al., 2019). Beck et al. (Beck et al., 2018) presented a case study of a DAO in Swarm City (City, 2023), a decentralized commerce platform. A recent work categorized the governance of several blockchains such as Bitcoin, Ethereum, Tezos, Polkadot, and some governance protocols like Uniswap (Adams et al., 2021), MakerDAO (MakerDAO, 2023) and Compound (Leshner and Hayes, 2019) into different types (Kiayias and Lazos, 2023). These invaluable prior work do not, nevertheless, empirically examine the data on existing DAOs to characterize how users interact with on-chain governance smart contracts.
7. Concluding discussion
In this work, we analyzed data from the Ethereum blockchain related to Compound and Uniswap, two widely used smart contracts. Our analysis centers on their decentralized governance protocols, with a particular focus on amendments to their contract. We found that the Compound contract is being actively amended—token holders continuously propose amendments that are then voted on by other token holders. We observed a striking concentration of tokens (be it in terms of their ownership or their delegation or their voting participation) in the hands of a few participants for both Compound and Uniswap, which raises serious concerns about the extent to which governance is decentralized in practice. For instance, our analysis showed that, on average, only $3.25$ voters were needed for the proposals to reach quorum and pass in Compound, and only $2.84$ voters were needed to reach $50$% of the total votes. Our analysis also highlights issues with the Compound use of on-chain voting – in particular, the transaction fees voters must pay to cast an on-chain vote can make it prohibitively expensive for voters with fewer tokens. These costs have implications for voting participation and can affect how voters, proposers, and other stakeholders interact with these protocols.
Our findings motivate new research directions; for example, we find empirical evidence in support of recent proposals to redefine voting power based on social rewards, such as a voter’s reputation or a voter’s contributions to the protocol (Sharma et al., 2023; Liu et al., 2022; Guidi et al., 2021) or the use of a quadratic voting scheme, where voting power is calculated as the square root of the number of tokens held by the voters (Buterin et al., 2019; Lalley and Weyl, 2018). We acknowledge that collecting data from smart contracts and determining wallet ownership can be a challenging task that requires the use of an Ethereum archive node. Therefore, we intend to make our data sets publicly available to enable scientific reproducibility of our results.
In conclusion, our research adds to the recent and growing body of knowledge on governance protocols and demonstrates the importance of measuring and auditing such systems to ensure they are working as expected. We hope that our study will be useful to future researchers in this field.
Acknowledgments
Appendix A Temporal dynamics of voting
@@
According to our analysis, voters take an average of $1.4$ days (with a std. of $0.95$ and a median of $1.34$ days) to cast their votes after the voting period began. The shortest and longest recorded delays in our data set are 00 and $3.39$ days, respectively. Fig. 18 shows the distribution of the time it takes voters to cast their votes for each proposal. We also highlight the voting delays for all votes cast per proposal in Fig. 21 in §D.
@@
Appendix B Filtering events to construct our Compound data set
@@
Appendix C Inferring Wallet Addresses Ownership
@@
Furthermore, to extend the available identified addresses in our analysis, we obtained the addresses of $2783$ identified users from the Sybil-List (Sybil, 2023b), a project maintained by Uniswap that uses cryptographic proofs to verify wallet addresses ownership. By combining the identified addresses from both sources, we were able to obtain the ownership of $3191$ inferred public wallet addresses to use in our analysis. We were able to infer $114$ ($3.41$%) out of the $3341$ unique addresses in our data set. Considering the top 10 most powerful voters for each proposal (refer to Fig. 20 in §D), we were able to infer $67$ ($50.37$%) of the $133$ unique addresses. Overall, our methodology allowed us to partially overcome the anonymity of public wallet addresses on the Ethereum blockchain and shed light on the ownership of these addresses in our data set. Finally, as an entity can control more than one address, we grouped the addresses we identified belonging to the same entity together to conduct our analysis.
Appendix D How voters cast their votes
This section examines how each of the top-10 voters of Compound and Uniswap cast their votes. Some proposals may not have received any votes if they were cancelled before the voting period began. See §4.2.3 for details. Fig. 20 shows how each of the top-10 voters cast their votes in each of the $126$ ($94.74$%) out of $133$ Compound proposals.